two parter! THANKS FOR THE HELP

1) A conducting sphere of radius 13.0 cm has a net charge of 2.2 x 10-8 C. What is the electric field at the surface of the sphere? Give your answer in N/C (equivalent to V/m).

2) A conducting sphere of radius 16.0 cm has a net charge of 2.8 x 10-8 C. If V=0 at infinite distance, what is the electric potential at the sphere's surface? Give your answer in volts.
Answer #1

Solution:

Question (1) Solution:

Given:

Net Charge (Q) = 2.20 x 10-8 C

and, r = 13.0 cm = 0.13 m

Therefore: the electric field at the surface of the sphere = kQ / r2 = {(9 x 109)(2.20 x 10-8)} / (0.13)2 = 11715.98 N/C

Question (2) Solution:

Given:

Net Charge (Q) = 2.8 x 10-8 C

and, r = 16.0 cm = 0.16 m

Therefore: the electric potential at the sphere's surface = kQ / r = {(9 x 109)(2.8 x 10-8)} / (0.16) = 1575 V

_____________________________________________________________________________________


Alright Dude, If that worked for you... don't forget to give THUMBS UP.(that will work for me!)
Please Vote...
If I missed something feel free to leave a comment.
atleast before giving down vote.
and, Thanks for using Chegg- Smarter way to study.